OpenStudy (anonymous):

An equilateral triangle has an altitude length of 36 feet. Determine the length of a side of the triangle.

OpenStudy (campbell_st):

well you need a drawing  the altitude bisects the base... so you can use pythagoras' theorem ... \[36^2 + (\frac{1}{2} x)^2 = x^2\] just solve for x

OpenStudy (anonymous):

@Campbell_st wait I got for the shortes side 12 square root 3? is that right?